# CAAP: Context-Aware Action Planning Prompting to Solve Computer Tasks with Front-End UI Only

### v1

@@ Abstract
Software robots have long been deployed in Robotic Process Automation (RPA) to automate mundane and repetitive computer tasks. The advent of Large Language Models (LLMs) with advanced reasoning capabilities has set the stage for these agents to now undertake more complex and even previously unseen tasks. However, the LLM-based automation techniques in recent literature frequently rely on HTML source codes for input, limiting their application to web environments. Moreover, the information contained in HTML codes is often inaccurate or incomplete, making the agent less reliable for practical applications. We propose an LLM-based agent that functions solely on the basis of screenshots for recognizing environments, while leveraging in-context learning to eliminate the need for collecting large datasets of human demonstration. Our strategy, named Context-Aware Action Planning (CAAP) prompting encourages the agent to meticulously review the context in various angles. Through our proposed methodology, we achieve a success rate of 94.4% on 67 types of MiniWoB++ problems, utilizing only 1.48 demonstrations per problem type. Our method offers the potential for broader applications, especially for tasks that require inter-application coordination on computers or smartphones, showcasing a significant advancement in the field of automation agents. Codes and models are accessible at https://github.com/caap-agent/caap-agent.
**footnotetext: These authors contributed equally to this work.
Keywords Large Language Model  $\cdot$ Autonomous Agent  $\cdot$ MiniWoB++
@@ 1 Introduction
Artificial intelligence (AI) agents are often commercialized as Robotic Process Automation (RPA) tools, designed to streamline business processes with minimal human intervention. Traditional RPA agents use rule-based algorithms to automate structured tasks, making them particularly effective for repetitive and standardized data processing. However, most desktop tasks are not well-suited to this rule-based approach, as they frequently involve handling unexpected situations and exceptions that are too numerous and varied to anticipate and program for in advance. In contrast, agents utilizing deep neural networks can adapt more flexibly to unforeseen situations. Studies have shown that agents trained on extensive data of expert behavior can effectively handle unseen tasks by mimicking the decision-making styles of human experts (Humphreys et al. (2022); Shaw et al. (2023)). The emergence of Large Language Models (LLMs) is now further advancing the capabilities of RPA agents. LLMs have demonstrated the potential to support complex decision-making processes with their advanced reasoning capabilities. Their In-Context Learning (ICL) ability allows them to learn meaningful actions for specific tasks with just a few examples, significantly reducing the need for large-scale data collection of expert demonstrations. As a result, research on agents leveraging ICL has been vigorously pursued in recent years (Yao et al. (2022a); Sridhar et al. (2023); Zheng et al. (2023); Shinn et al. (2024); Lin et al. (2023); Huang et al. (2022); Liu et al. (2024)). As AI chips are increasingly integrated into computers and smartphones (Intel (March 26, 2024); Samsung (January 18, 2024)), and as these AI devices are poised to become the standards in the market, this new trend in the design of AI-driven RPA tools is likely to have a profound impact.
@@
One of the key challenges in developing reliable LLM-based agents is ensuring their accurate interpretation and representation of real-world information. This task, commonly referred to as grounding (Liu (2023)), involves various aspects such as accurately detecting UI objects on the screen, identifying appropriate actions based on the context, and removing ambiguity in communication through a structured, predefined format. Complications arising from grounding issues can be mitigated by employing software-level shortcuts, such as using HTML or the Document Object Model (DOM) as input sources, or interacting with the digital environment through well-designed application programming interfaces (APIs) (Kim et al. (2023); Sun et al. (2023)). However, the effectiveness of these methods depends on the availability of such shortcuts. They lack general applicability to broader desktop tasks that are not web-based or that require the use of applications with no provided APIs. Often, many applications and web services in the real world do not maintain rigorous development standards.
@@
In this paper, we introduce an agent model that exclusively utilizes visual information from the screen as input. This approach is intuitive, considering that computer-based tasks are inherently designed for human operators who primarily perceive and interpret their environments through visual stimuli on the computer screen. Our agent emulates this human perceptual process. To improve the agent’s ability to process this input and enhance its effectiveness in task completion, we have developed a novel prompting technique named Context-Aware Action Planning (CAAP). This technique involves systematically organizing the contextual information necessary for making action decisions and identifying syntactic structures that induce the most effective Chain of Thought (CoT) (Wei et al. (2022)). This enables the agent to reason about its environment and autonomously generate information essential for decision-making. Our agent demonstrates superior performance on the MiniWoB++ benchmark, achieving high task-solving accuracy across a diverse array of tasks. It surpasses previous studies by utilizing only a minimal amount of human demonstration data, thereby illustrating its efficiency and effectiveness in handling complex computer tasks with limited training cost.
@@
•  We introduce a novel design for a software robot (agent) that operates exclusively through human-oriented front-end UI channels for both input and output. This approach eliminates the constraints associated with HTML/DOM and application-specific API calls. Our LLM-based agent leverages ICL, allowing it to adapt to new tasks effortlessly without requiring extensive neural network retraining.
•  We introduce an effective prompting technique named CAAP that enhances the ICL ability of an LLM-based agent in managing complex desktop tasks. By systematically structuring contextual information and leveraging syntactic patterns that trigger optimal CoT reasoning, CAAP significantly outperforms other agent designs.
•  We propose fairer metrics for comparing the performance of agents in the MiniWoB++ benchmark. Prior to our work, each study reported results on a selectively chosen subgroup of tasks, making direct comparisons challenging.
@@ 2 Related work
Vision models for agents.
Numerous advanced techniques have been developed for detecting and interpreting UI elements on screens. Faster R-CNN (Ren et al. (2015)) employs a Region Proposal Network (RPN) to extract areas of interest, followed by accurate object detection within these regions. This approach offers high accuracy in complex scenes, though at the cost of slower processing speed. The Single Shot MultiBox Detector (SSD) (Liu et al. (2016)) offers rapid inference by eliminating the RPN and maintains decent accuracy. To improve small object detection and reduce false positives, one can diversify default boxes and feature maps for the SSD. Co-DETR (Zong et al. (2023)), which leverages a Transformer-based architecture, performs detection while learning the relationships between objects. This anchor-free approach simplifies the learning process and effectively understands object interrelations, but it requires significant time and resources for training. YOLOv8 (Redmon et al. (2016)) processes predictions for object classes and bounding boxes across the entire image in one go. YOLOv8 has both high speed and high accuracy compared to most other models. Unlike the previously mentioned technologies that primarily focus on extracting positional and typological information of objects, Pix2Struct (Lee et al. (2023)) specializes in understanding detected objects. Designed around a Vision Transformer (ViT) (Dosovitskiy et al. (2020)) and a T5 transformer (Raffel et al. (2020)), it takes an image as input and outputs structured text, playing a crucial role in extracting and analyzing detailed attributes and embedded text information. Our approach leverages the combination of YOLOv8 and Pix2Struct for visual information processing, aiming to capitalize on the strengths of both high-speed object detection and detailed attribute analysis.
Deep learning approaches
Nakano et al. (2021) propose WebGPT, a GPT-3 model fine-tuned using both supervised learning (SL) and reinforcement learning (RL) on a human feedback dataset. Similarly,  Furuta et al. (2023) present WebGUM, which combines Vision Transformer (ViT) and T5 transformers, trained via SL on a large-scale demonstration set of 401K examples. Both models utilize HTML inputs to tackle computer-based tasks. Humphreys et al. (2022) introduce CC-Net, a multimodal model that integrates ResNet and transformers. Designed to predict the next action from screenshots and Document Object Model (DOM) inputs, CC-Net relies heavily on DOM and requires both SL and RL for optimal performance. Conversely,  Shaw et al. (2023) propose the Pix2Act model, based on Pix2Struct (Lee et al. (2023)), which is the first agent model to achieve high performance solely from screenshots, also trained using both SL and RL. While these methods excel in their respective domains, they face common challenges. Each requires large-scale expert task-solving demonstrations, complicating adaptation to new tasks. Additionally, the integration of RL results in more complex training procedures.
LLM-based agents with ICL.
Yao et al. (2022a) develop the ReAct model, which assesses the current situation before determining the subsequent action.  Kim et al. (2023) introduce RCI prompting, which enhances action planning through progressive self-critique, underscoring the vital role of advanced reasoning techniques in agent development.  Huang et al. (2022) suggest a mechanism known as Inner Monologue (IM), actively incorporating real-time environmental states into action planning via an explicit feedback loop. Inspired by Inner Monologue,  Sun et al. (2023) create AdaPlanner, prompting LLMs to generate action plans as Python functions and adjust these plans should the execution fail. Both IM and AdaPlanner emphasize the importance of continuous adaptation to environmental changes for the effective completion of complex tasks.  Zheng et al. (2023) introduce the Synapse framework, leveraging successful past action trajectories as few-shot exemplars to guide future actions. Similarly,  Shinn et al. (2024) present Reflexion, a method that encourages LLMs to reflect on past failures, document these incidents, and utilize the acquired insights in subsequent tasks.
Grounding techniques
Zheng et al. (2023) demonstrate that agents with human involvement in grounding tasks significantly outperform those entirely driven by LLMs, underscoring the importance of acquiring sophisticated grounding techniques for language models to effectively function as agents.  Baechler et al. (2024) introduce ScreenAI, a model that addresses visual understanding by answering queries about the visual state depicted in screenshots, such as "What text is written in the search bar?" ScreenAI leverages the object understanding capabilities of Pix2Struct and the instruction understanding abilities of T5 to approach visual grounding issues.  Yang et al. (2023) propose the Set-of-Mark by enhancing screenshots with numerical labels via MaskDINO (Li et al. (2023)) and integrating these annotated screenshots with text instructions delivered to GPT-4V to tackle visual grounding challenges. He et al. (2024) expand Set-of-Mark into the agent domain, combining it with ReAct prompting in the WebVoyager agent model.  Cheng et al. (2024) propose SeeClick, which uses only screenshots as input and predicts the next click-based action by dividing the screenshot into a grid and training the model to learn grid coordinates to determine click actions on target objects.
@@
3 Front-end UI agents for computer tasks
Our CAAP-based agent is engineered to replicate the human approach to solving computer tasks. Typically, humans manage complex tasks by first assessing the current situation through visual observation of the monitor and interpreting the displayed information. The complex task is broken down into smaller, manageable actions. Based on the observed state, they select the most appropriate action and execute it using a keyboard and mouse. Following this, they review the updated state on the screen and cyclically repeat the processes of perception, interpretation, decision-making, and execution until the task is completed. To mirror this human-like process, we have designed our agent with three key components: Visual Observer [^1] Notes on the possible use of a large multimodal model (LMM): We expect that in the near future, large multimodal models (LMMs) will become advanced enough to replace our current Visual Observer. By using an LMM instead of the Visual Observer, our agent design will be more adaptable to various practical applications because training neural networks will no longer be necessary. Instead, we will rely solely on the integrated capabilities of the LMM for both observing (Visual Observer) and decision-making (Decision Maker). We also believe that our CAAP prompting technique, which uses a modular approach to separate sensing and reasoning, will remain superior to an end-to-end system where a single LMM handles both observation and decision-making tasks. This is because our approach more closely mirrors how humans interact with computer tasks., Decision Maker, and Action Executor. This structure is illustrated in Figure 1.
3.1 Visual observer
To convert visual information from a screen into the linguistic domain, we employ a two-step approach. Initially, in the detection stage, UI elements are identified and their spatial locations are extracted. Subsequently, these elements undergo further analysis through advanced object understanding techniques, during which detailed attributes of each element are captured. This processing results in the information being organized into structured text, ready to be handed to Decision Maker as the representation of the current state.
@@ 3.1 Visual observer
For the detection and extraction of location information of all UI elements present in screenshots, we fine-tuned the YOLOv8-x model. The YOLO model, known for its rapid and precise detection in one shot, typically extracts object type information along with bounding boxes. However, we utilize only the bounding boxes, leaving the extraction of type information to the subsequent UI element understanding module for greater accuracy.
UI element understanding
We use the Pix2Struct model (Lee et al. (2023)) to extract detailed attributes from all UI elements identified during the detection stage. We fine-tune the original Pix2Struct model to match the specific output format required by our target environment. During this fine-tuning process, we employ a different masking technique from Pix2Struct style, as depicted in Figure 2 (a). Our masking technique highlights the target UI element by outlining its region and darkening the surrounding image areas, as demonstrated in Figure 2 (b). This method preserves the pixel-level details of the targeted element while helping the model learn the spatial relationships between the element and its adjacent elements. The structured output from the model is presented as shown in Figure 2 (c), with the range of extractable attributes for each element detailed in Table 2 of Appendix B.
@@ 3.2 Decision maker
Upon receiving the representation of the current screen, Decision Maker module is employed to determine possible actions for completing the task. The decision-making process is facilitated by its two key components: a context collector and a CAAP prompter. The context collector gathers vital contextual information that aids in task resolution. This includes exemplary demonstrations, the ultimate task goal, the history of actions performed so far, and the range of possible actions that the language model can choose from, all formatted as text. The CAAP prompter then uses this information to construct prompts that are sent to the language model, eliciting responses that guide the next steps. The CAAP prompting method, one of our major contributions in this paper, is discussed in detail in Section 4.
@@ 3.3 Action executor
The actions determined by the decision maker are implemented through a low-level interface to effect changes in the environment by the Action Executor module. It first identifies a series of recommended actions by parsing the response text from the Decision Maker. Then, it executes these actions using the keyboard and mouse interfaces of the operating system. After each action, it receives an updated screenshot, halting further actions if a change in the environment is detected. The types of actions that can be executed include mouse actions such as clicking, pointing, and scrolling, as well as keyboard actions like typing and using common shortcuts. These are enumerated in Table 3 of Appendix C.

### v2

@@ Abstract
Software robots have long been used in Robotic Process Automation (RPA) to automate mundane and repetitive computer tasks. With the advent of Large Language Models (LLMs) and their advanced reasoning capabilities, these agents are now able to handle more complex or previously unseen tasks. However, LLM-based automation techniques in recent literature frequently rely on HTML source code for input or application-specific API calls for actions, limiting their applicability to specific environments. We propose an LLM-based agent that mimics human behavior in solving computer tasks. It perceives its environment solely through screenshot images, which are then converted into text for an LLM to process. By leveraging the reasoning capability of the LLM, we eliminate the need for large-scale human demonstration data typically required for model training. The agent only executes keyboard and mouse operations on Graphical User Interface (GUI), removing the need for pre-provided APIs to function. To further enhance the agent’s performance in this setting, we propose a novel prompting strategy called Context-Aware Action Planning (CAAP) prompting, which enables the agent to thoroughly examine the task context from multiple perspectives. Our agent achieves an average success rate of 94.5% on MiniWoB++ and an average task score of 62.3 on WebShop, outperforming all previous studies of agents that rely solely on screen images. This method demonstrates potential for broader applications, particularly for tasks requiring coordination across multiple applications on desktops or smartphones, marking a significant advancement in the field of automation agents. Codes and models are accessible at https://github.com/caap-agent/caap-agent.
@@ 1 Introduction
Artificial intelligence (AI) agents are often commercialized as Robotic Process Automation (RPA) tools, designed to streamline business processes with minimal human intervention. Traditional RPA agents use rule-based algorithms to automate structured tasks, making them particularly effective for repetitive and standardized data processing. However, most desktop tasks are not well-suited to this rule-based approach, as they frequently involve handling unexpected situations and exceptions that are too numerous and varied to anticipate and program for in advance. In contrast, agents utilizing deep neural networks can adapt more flexibly to unforeseen situations. Studies have shown that agents trained on extensive data of expert demonstrations can effectively handle unseen tasks by imitating the decision-making styles of human experts (Humphreys et al., 2022; Shaw et al., 2023). The emergence of Large Language Models (LLMs) is now further advancing the capabilities of RPA agents. LLMs have demonstrated the potential to support complex decision-making processes with their advanced reasoning capabilities. Their In-Context Learning (ICL) ability allows them to learn meaningful actions for specific tasks with just a few examples, significantly reducing the need for large-scale data collection of expert demonstrations. As a result, research on agents leveraging ICL has been vigorously pursued in recent years (Yao et al., 2022a; Sridhar et al., 2023; Zheng et al., 2023; Shinn et al., 2024; Lin et al., 2023; Huang et al., 2022; Liu et al., 2024). As AI chips are increasingly integrated into computers and smartphones (Intel, March 26, 2024; Samsung, January 18, 2024), and as these AI devices are poised to become the standards in the market, this new trend in the design of AI-driven RPA tools is likely to have a profound impact.
@@
Most prior research on agents for solving computer tasks has primarily relied on HTML or the Document Object Model (DOM) as input sources, or has leveraged well-structured Application Programming Interfaces (APIs) to interact with digital environments (Humphreys et al., 2022; Liu et al., 2018; Gur et al., 2022; Kim et al., 2023; Sun et al., 2023; Furuta et al., 2023). However, the effectiveness of these methods depends on the availability of such shortcuts. They lack general applicability to broader desktop tasks that are not web-based or that require the use of applications with no provided APIs. Often, many applications and web services in the real world do not maintain rigorous development standards.
@@
In this paper, we present an agent model that exclusively uses visual information from the screen as input and keyboard/mouse actions as output. This approach is intuitive, given that many computer-based tasks are designed specifically for human operators. Our proposed model features a modular architecture that distinctly separates the vision model, responsible for perceiving the environment, from the decision-making model, which handles action planning. This clear separation not only enhances the agent’s functionality but also improves its adaptability across various applications. To further support and maximize the effectiveness of our agent design, we have developed a novel prompting technique called Context-Aware Action Planning (CAAP). This technique systematically organizes the contextual information required for action decisions and identifies syntactic structures that trigger the most effective Chain of Thought (CoT) (Wei et al., 2022). Our agent demonstrates superior task-solving performance on the MiniWoB++ and WebShop benchmark. It surpasses previous studies by utilizing only a minimal amount of human demonstration data, thereby illustrating its efficiency and effectiveness in handling complex computer tasks with limited training cost.
@@
•  We present the first LLM-based agent for computer tasks that leverages ICL while emulating human behavior by using front-end User Interface (UI) channels as input-output sources. It processes input through an image-to-text model followed by an LLM, a combination that eliminates the need for extensive human labor in dataset collection before deploying the agent for new tasks. The agent operates exclusively through keyboard and mouse actions, enabling deployment on any GUI-based system without the need for special APIs to be provided.
•  We introduce an effective prompting technique named CAAP, which enhances the ICL capabilities of an LLM-based agent in managing complex desktop tasks. By systematically structuring contextual information and leveraging syntactic patterns that trigger optimal CoT reasoning, CAAP significantly outperforms other agent designs.
@@ 2 Related work
Agents with image inputs
Humphreys et al. (2022) introduced CC-Net, a compound model of ResNet, Transformer, and LSTM that uses image and DOM data as inputs and produces policies for keyboard-mouse actions. It was trained via Supervised Learning (SL) and Reinforcement Learning (RL), covering the largest number of MiniWoB++ tasks to date (104 tasks). Furuta et al. (2023) proposed WebGUM, which combines Vision Transformer (ViT) and T5 transformers. It was trained using SL to solve both MiniWoB++ and WebShop tasks, utilizing both image and HTML inputs. Shaw et al. (2023) introduced Pix2Act, the first agent to completely eliminate reliance on DOM or HTML for solving MiniWoB++ and WebShop tasks. The Pix2Act model is built upon the Pix2Struct architecture (Lee et al., 2023) and uses only image inputs. While these studies provide valuable insights and serve as important benchmarks, their methods are impractical for real-world applications as they require around a million or more human demonstration data samples, demanding thousands of hours of human labor for training. Cheng et al. (2024) introduced SeeClick, a method based on a Large Multimodal Model (LMM). Aided by the strong reasoning capabilities of the LMM, SeeClick required a much smaller dataset of 2.8K human demonstrations to train for MiniWoB++ tasks. It fell short in overall performance, however, achieving less than a 70% success rate on the 55 tasks it covered. A common feature of all these approaches is the integration of image interpretation and action planning models within a single architecture. This rigid design poses a significant limitation: when new UI elements need to be recognized or additional actions supported for a new task, the entire architecture must be retrained.
LLM-based agents with ICL
To reduce dependence on large quantities of human demonstration data, several studies have explored the use of LLMs for action planning. Yao et al. (2022a) developed the ReAct model, which evaluates the current state before determining subsequent actions. Kim et al. (2023) introduced the Recursive Criticism and Improvement (RCI) agent, which enhances action planning through iterative self-critique, highlighting the importance of advanced reasoning techniques in agents. Huang et al. (2022) proposed Inner Monologue (IM), a mechanism that integrates real-time environmental states into action planning via an explicit feedback loop. Inspired by IM, Sun et al. (2023) introduced AdaPlanner, which prompts LLMs to generate action plans as Python functions and refine them based on feedback from execution failures. Despite their innovations, all these approaches share a common limitation: their architectures are designed to operate solely with HTML inputs, restricting their applicability to constrained, well-curated web environments.
3 Modularized GUI agent for computer tasks
@@
Our agent consists of three distinct modules: a visual observer that processes and interprets image inputs, a decision-maker responsible for action planning, and an action executor that carries out the planned actions. The agent operates by repeatedly executing a cycle consisting of a series of sequential processes involving these three modules. Each of these modules is further divided into functional submodules, as depicted in Figure 1. This approach offers several advantages, including modular updates, better scalability, and parallelized processing:
Modular updates
The modular architecture provides the flexibility to perform isolated updates, minimizing disruption to the overall system and allowing for more efficient model improvements. In contrast, end-to-end monolithic agent designs, where perception and decision-making are tightly integrated within one LMM, require updates to the entire model, which is time-consuming and resource-intensive.
Scalability
The agent’s modular structure supports horizontal scalability by allowing the addition of new functional submodules to enhance overall performance. For instance, a reflection submodule (Shinn et al., 2024; Zhao et al., 2024) could be added to the decision-maker module, enabling the agent to analyze and learn from previous actions or experience. The visual observer module could be supplemented with multiple vision models, each specialized for different domains or tasks. This allows the system to dynamically select and apply the appropriate model based on the context, significantly enhancing the agent’s adaptability to various environments.
Parallelization
The agent’s modular design also supports parallel processing of image models. For example, the screen image can be divided into multiple sections, with multiple instances of visual observers processing different regions concurrently. This parallel processing can be integrated seamlessly without disrupting the performance of the decision-making modules.
3.1 Visual observer***These authors contributed equally to this work.
The visual observer module employs a two-step approach to transform visual information from a screen into the linguistic domain. Initially, in the detection stage, UI elements are identified, and their spatial locations are extracted. Subsequently, these elements undergo further analysis through advanced object understanding techniques, during which detailed attributes of each element are captured. This processing results in the information being organized into structured text, ready to be handed to the decision-maker as the representation of the current state.
**footnotetext: Notes on the potential use of an LMM: We anticipate that in the near future, LMMs may advance enough to replace our current visual observer. The adoption of LMMs would reduce the need for training vision models tailored to specific domains, further simplifying the application of our agent design across various practical use cases. However, our CAAP prompting technique, which separates sensing and reasoning into distinct modules, will continue to be superior to end-to-end systems where a single LMM manages both observation and decision-making tasks. This modular approach will maintain significant advantages, including easier updates, improved scalability, and enhanced parallelization.
@@
For the detection and extraction of location information of all UI elements present in screenshots, we fine-tuned the YOLO model. The YOLO model, known for its rapid and precise detection in one shot, typically extracts object type information along with bounding boxes. However, we utilize only the bounding boxes, leaving the extraction of type information to the subsequent UI element understanding module for greater accuracy.
@@
UI element understanding
We use the Pix2Struct model (Lee et al., 2023) to extract detailed attributes from all UI elements identified during the detection stage. We fine-tune the original Pix2Struct model to match the specific output format required for our task. During this fine-tuning process, we employ a different masking technique from Pix2Struct style, as depicted in Figure 2 (a). Our masking technique highlights the target UI element by outlining its region and darkening the surrounding image areas, as demonstrated in Figure 2 (b). This method preserves the pixel-level details of the targeted element while helping the model learn the spatial relationships between the element and its adjacent elements. The structured output from the model is presented as shown in Figure 2 (c), with the range of extractable attributes for each element detailed in Table 7 of Appendix C.
@@ 3.2 Decision maker
Upon receiving the representation of the current screen, the decision maker module is employed to determine possible actions for completing the task. The decision-making process is facilitated by its two key components: a context collector and a CAAP prompter. The context collector gathers vital contextual information that aids in task resolution. This includes exemplary demonstrations, the ultimate task goal, the history of actions performed so far, and the range of possible actions that the language model can choose from, all formatted as text. The CAAP prompter then uses this information to construct prompts that are sent to the language model, eliciting responses that guide the next steps. The CAAP prompting method, one of our major contributions in this paper, is discussed in detail in Section 4.
@@ 3.3 Action executor
The actions determined by the decision maker are implemented through a low-level interface to effect changes in the environment by the last module, action executor. It first identifies a series of recommended actions by parsing the response text from the decision maker. Then, it executes these actions using the keyboard and mouse interfaces of the operating system. After each action, it receives an updated screenshot, halting further actions if a change in the environment is detected. The types of actions that can be executed include mouse actions such as clicking, pointing, and scrolling, as well as keyboard actions like typing and using common shortcuts. These are enumerated in Table 8 of Appendix D and implemented as Python functions using the PyAutoGUI package.